OpenStudy (anonymous):

I will give a metal to whoever helps me. I would really appreciate it. What is the value of the x variable in the solution to the following system of equations? 7x - 2y = 21 4x + y = 57 21 -21 9 -9

OpenStudy (akashdeepdeb):

Now use it in the second equation! y = 57-4x [Okay?] Put it in the first equation as 7x - 2y = 21 7x - 2(57-4x) = 21 Solve the equation now! :) Understood? :)

OpenStudy (anonymous):

one sec. gonna solve it

OpenStudy (akashdeepdeb):

:D

OpenStudy (anonymous):

x=9 ^_^

OpenStudy (akashdeepdeb):

:)

OpenStudy (anonymous):

Thank you! :) I told you i only needed a refresher! LOL

OpenStudy (akashdeepdeb):

Hahah...Yeah!! :D